# Quasi-Monte Carlo with one categorical variable

### v2

@@ 1 Introduction
We study randomized quasi-Monte Carlo (RQMC) integration in a setting where one of the input variables is categorical, taking only some finite number $L\geqslant 2$ of different values. The main motivation for this problem is mixture sampling where the variables of interest come from $L$ different distributions representing $L$ different scenarios. That is, $\bm{x}\sim P_{\ell}$ with probability $\alpha_{\ell}>0$ where $\sum_{\ell=1}^{L}\alpha_{\ell}=1$. The distributions $P_{\ell}$ are commonly called mixture components. We call them ‘strata’ to avoid confusion with vector components. We want to estimate $\mu={\mathbb{E}}(g(\bm{x}))=\sum_{\ell=1}^{L}\alpha_{\ell}{\mathbb{E}}(g(\bm{x})\!\mid\!\bm{x}\sim P_{\ell})$. Our approach will also allow us to use distinct functions $g_{\ell}$ within each stratum.
A straightforward approach to this problem is to simply replace a standard Monte Carlo (MC) workflow with RQMC sampling. In such a workflow, one random variable selects stratum $\ell$ with probability $\alpha_{\ell}$. Then subsequent random variables are used to generate $\bm{x}\sim P_{\ell}$. When MC is replaced by RQMC, we find some advantages to sampling the $L$ strata more equally than their nominal proportions $\alpha_{\ell}$ would indicate.
@@
An alternative to the standard MC workflow is to compute $L$ different independently sampled RQMC estimates, one per stratum, and combine them according to their mixture probabilities $\alpha_{\ell}$. That approach misses a potential variance reduction from RQMC sampling when the within-stratum estimates are negatively correlated. Such negative correlation need not happen; we give a reason to expect that at the end of the paper.
for $\bm{x}_{i}\sim\sum_{\ell=1}^{L}\alpha_{\ell}p_{\ell}$. By Jensen’s inequality, this has lower Monte Carlo variance than we would get from
$\sum_{\ell=1}^{L}\frac{\alpha_{\ell}}{n_{\ell}}\sum_{i=1}^{n_{\ell}}\frac{g(\bm{x}_{i,\ell})p(\bm{x}_{i,\ell})}{p_{\ell}(\bm{x}_{i,\ell})}.$
working separately within each of the $L$ strata.
@@
RQMC does best under favorable regularity assumptions, such as continuity, on the integrand. The presence of the categorical variable makes the resulting integrand discontinuous. That discontinuity is however axis parallel and so it is a QMC-friendly discontinuity in the terminology of Wang and Tan (2013). In some very regular cases, RQMC can attain a mean squared error (MSE) of $O(n^{-3+\epsilon})$. Such a rate can only be attained over a sequence of special sample sizes that grow at least exponentially (Sobol’, 1998; Owen, 2016). This rate can be attained by scrambled Sobol’ points with sample size equal to a power of $2$. Even an integrand with a QMC-friendly axis-parallel discrepancy is not generally integrated with an error of this rate. By using an importance sampling (Owen, 2013, Chapter 9) adjustment, we can arrange to obtain $n_{\ell}=2^{m_{\ell}}$ points from $P_{\ell}$ while also having $n=\sum_{\ell=1}^{L}n_{\ell}$ be a power of $2$. This can then give the $O(n_{\ell}^{-3+\epsilon})$ rate within every stratum and hence overall.
@@
This paper is organized as follows. Section 2 introduces our notation, gives some background on RQMC, and proves some results about within-stratum discrepancies. Section 3 describes how to allocate samples to strata. We see that in RQMC sampling, it is advantageous to use more nearly equal allocations than those given by the stratum sizes. Section 4 shows how to partition a sample size $n=2^{m}$ into $L$ within-stratum sample sizes $n_{\ell}=2^{m_{\ell}}$. Section 5 establishes some minimax variance results from using equal or nearly equal sample sizes for the strata. Section 6 shows a simple toy example with $L=8$ strata where RQMC with $n_{\ell}$ taken to be powers of two does best. Section 7 uses a mixture model version of the Saint-Venant flood example from Limbourg and De Rocquigny (2010) where non-nominal sampling proportions do best. Section 8 has some discussion and conclusions. The proof of one of our results is in Appendix A.
@@
We finish this introduction by describing some related work in the literature. Our discussion uses the notation of this paper to describe the findings of the other papers. The key differences between our work and the prior work we found is in how the problem has been formulated, and how the values are sampled.
@@
We use the first component of our RQMC points to select the stratum to sample from. Keller and Grünschloß (2012) used that approach in parallel computation over $L$ processors. Their first component was used to determine which processor $\ell\in\{1,2,\dots,L\}$ got the $i$’th job. The remaining components of their QMC points sampled that job on processor $\ell$. In that setting, $\alpha_{\ell}=1/L$ and it makes sense to choose $L$ to be a power of $2$ at least for Sobol’ sequences.
@@
Cui et al. (2025) consider a target distribution on a bounded hyperrectangle $[\bm{a},\bm{b}]\subset^{d}$. The target probability density function $\pi$ is bounded. They approximate it by a sum of products of hat functions. That is equivalent to a mixture of distributions whose components have independent triangle distributions. They consider adaptive methods to select those triangle densities given an unnormalized $\bar{\pi}\propto\pi$. For a possibly large number of mixture components, they sample each component with $n_{j}\doteq n\alpha_{j}$ input values. They use the first $n_{j}$ points of the same QMC point set for each of the strata. Their integrands $f$ belong to an $L_{p}$ weighted space with a slightly different norm from the usual ones. Their equation (3) defining the norm also uses some sup-norms of absolute mixed partial derivatives. Then sampling on a class of digital nets, they are able to bound the integration error by an expression using the integrand’s norm. They have another bound when the density satisfies a Hölder condition.
@@
We use different RQMC points in each stratum. Our inputs are randomized inputs, and so we expect better error cancellation than we would get using the same inputs for every stratum. Since our analysis is based on variances, we do not require a bounded domain or integrand. We also find an advantage, theoretically and empirically, to diverging from proportional allocation.
@@
Klebanov and Sullivan (2023) describe using transport maps from a $\mathcal{N}(0,I)$ reference distribution to sample a mixture like $P=\sum_{\ell=1}^{L}\alpha_{j}\mathcal{N}(\mu_{j},\Sigma_{j})$. They construct a continuous mapping over $t\in[0,1]$ with $\mathcal{N}(0,I)$ at $t=0$ reaching $P$ at $t=1$ and solve an ordinary differential equation along that path. For $\ell<L$, they take the first $n_{\ell}=\lfloor n\alpha_{\ell}\rfloor$ points of a quasi-Monte Carlo sequence and use those to sample from component $\ell$. Then $n_{L}=n-\sum_{\ell=1}^{L-1}n_{\ell}$. This is different from our strategy of using the first RQMC component to select the stratum which we prefer because it avoids giving two within-stratum samplers the same points. They do not require $n\gg L$. Some of our expressions only require $n>2L$, but as we use asymptotic variances within strata, our results are most relevant for large $n/L$. They consider more general location-scale families than the Gaussian distribution and they can even include some $\alpha_{\ell}<0$ so long as $P$ never has a negative probability density function.
@@ 2 Background
Here we introduce notation to describe the different variables used in the integration workflow. Then we define digital nets and discuss their randomizations. For a broader survey of RQMC see L’Ecuyer and Lemieux (2002).
We use $\mathbb{N}$ for the set of positive integers. We use $\bm{1}\{A\}$ for the indicator function that equals $1$ when $A$ holds and is 00 otherwise.
@@ 2.1 Integration variables
Our goal is to estimate
$\mu=\sum_{\ell=1}^{L}\alpha_{\ell}\int_{{}^{d}}g(\bm{x})p_{\ell}(\bm{x})\mathrm{\,d}\bm{x}$
@@
We will use RQMC points $\bm{z}\in[0,1]^{s+1}$. We partition $\bm{z}$ into its first component $v$ that we use to select a stratum $\ell$ and the subsequent $s$ components $\bm{u}$ that we use to sample $\bm{x}\sim P_{\ell}$. That is, $\bm{z}=(v,u_{1},u_{2},\dots,u_{s})$. For a function $\bm{\ell}:[0,1]\to\{1,2,\dots,L\}$, we will sample the $i$’th point using sample stratum $\bm{\ell}(v_{i})$. The pre-image of each $\ell\in\{1,2,\dots,L\}$ under $\bm{\ell}$ will always be a sub-interval of $[0,1]$. We let $\beta_{\ell}$ be the length of that subinterval. We abbreviate $\bm{\ell}(v_{i})$ to $\ell(i)$ in some expressions. Having selected stratum $\ell(i)$, we draw $\bm{x}_{i}\in^{d}$ via $\bm{x}_{i}=\phi_{\ell(i)}(\bm{u}_{i})$.
@@
We also define the within-stratum integrands individually as functions of $[0,1]^{s}$. The within-stratum integrands and integrals are
@@
At this point we can introduce a small generalization. Since we already have a dependence on $\ell$ in $h_{\ell}$ we could also replace $h_{\ell}(\bm{u})=g(\phi_{\ell}(\bm{u}))$ by $h_{\ell}(\bm{u})=g_{\ell}(\phi_{\ell}(\bm{u}))=g_{\ell}(\bm{x})$ to allow the function being applied to $\bm{x}$ to vary with $\ell$.
@@ 2.2 Sampling the strata
While $v=1$ has probability zero, for completeness we set $\bm{\ell}(v)=L$ when $v=1$. Because $v\sim\mathbf{U}[0,1]$, the expected number of samples from stratum $\ell$ is $n\beta_{\ell}$. We will always order the strata so that $\beta_{1}\geqslant\beta_{2}\geqslant\cdots\geqslant\beta_{L}$. That brings an advantage in Section 2.5.
We will assume that the sampling values $v_{1},\dots,v_{n}$ that we use are ‘stratified’. By this we mean that every interval $I_{k}=[(k-1)/n,k/n)$ contains precisely one of $v_{1},\dots,v_{n}$ for $k=1,\dots,n$. The next proposition shows that if every $\beta_{\ell}>2/n$, then every $n_{\ell}>0$ when $v_{i}$ are stratified.
@@
For $n\geqslant 1$, let $v_{1},\dots,v_{n}$ be stratified. For $[A,B)\subset[0,1)$, let $n_{*}$ be the number of $v_{i}\in[A,B)$. Then
@@
Write $A=r/n+a$ and $B=s/n+b$ for integers $0\leqslant r\leqslant s<n$ and $a,b\in[0,1/n)$. Then $\beta=t/n+c$ for $t=s-r$ and $-1/n<c<1/n$. If $r=s$, then $0\leqslant n_{*}\leqslant 1$ and then (6) holds trivially.
For $s>r$ we may write
$[A,B)=[A,(r+1)/n)\,\cup\,[(r+1)/n,s/n)\,\cup\,[s/n,s/n+b).$
Then $[A,B)$ contains at least $s-r-1$ of the intervals $I_{k}=[(k-1)/n,k/n)$ for $k=1,\dots,n$ and it intersects with at most $s-r+1$ of them. Therefore $|n_{*}-(s-r)|\leqslant 1$. Now $\lceil n\beta\rceil=s-r+\lceil nc\rceil$ and so $s-r\geqslant\lceil n\beta\rceil-1$ establishing the lower limit in (6). Similarly $\lfloor n\beta\rfloor=s-r+\lfloor nc\rfloor$ so $s-r\leqslant\lfloor n\beta\rfloor+1$ establishing the upper limit in (6). ∎
@@ 2.3 Local and star discrepancy
For $\bm{a}\in[0,1)^{d}$, the anchored box $[\bm{0},\bm{a})$ has volume $\mathrm{Vol}([\bm{0},\bm{a}))=\prod_{j=1}^{d}a_{j}$. The local discrepancy of $n$ points $\bm{x}_{1},\dots,\bm{x}_{n}\in[0,1)^{d}$ at $\bm{a}$ is
@@
The Koksma-Hlawka inequality (Hickernell, 2014) is that
@@ 2.4 Digital nets
Here we present the defining properties of digital nets. They exist in integer bases $b\geqslant 2$ (Dick and Pillichshammer, 2010) but we only consider base $2$ here. For a dimension $d\geqslant 1$, elementary intervals in base $2$ are dyadic hyperrectangles of the form
@@
for integers $k_{j}\geqslant 0$ and $0\leqslant c_{j}<2^{k_{j}}$. We will use $|\bm{k}|=\sum_{j=1}^{d}k_{j}$, so $H(\bm{c},\bm{k})$ has volume $2^{-|\bm{k}|}$.
For integers $m\geqslant t\geqslant 0$ the $n=b^{m}$ points $\bm{z}_{1},\dots,\bm{z}_{n}\in[0,1)^{d}$ are a $(t,m,d)$-net in base $b$ if $H(\bm{c},\bm{k})$ contains precisely $nb^{-|\bm{k}|}=b^{m-|\bm{k}|}$ of them, whenever $m-|\bm{k}|\geqslant t$. Smaller $t$ are better. We always have $t\leqslant m$ because $t=m$ just means that all the points are in $[0,1)^{d}$.
A $(t,d)$-sequence in base $2$ is an infinite sequence of values $\bm{x}_{i}\in[0,1]^{d}$, with the following property: any subsequence $\bm{x}_{a2^{r}+1},\dots,\bm{x}_{(a+1)2^{r}}$ for integers $a\geqslant 0$ and $r\geqslant m$ is a $(t,r,d)$-sequence in base $2$. With a $(t,d)$-sequence we get an infinite sequence of $(t,m,d)$-nets. Consecutive pairs of those $(t,m,d)$-nets yield an infinite sequence of $(t,m+1,d)$-nets of which pairs form $(t,m+2,d)$-nets and so on without limit. We will use digital nets in base $2$ taken from Sobol’ sequences using the direction numbers from Joe and Kuo (2008).
We call a $(t,m,d)$-net in base $2$ stratified if for each component $j=1,\dots,d$, the values $x_{1j},\dots,x_{nj}$ are stratified. A net with $t=0$ is automatically stratified, but some digital nets with $t>0$ are also stratified. The $(t,m,d)$-nets from Sobol’ sequences are stratified. The stratification property means we could use the $j$’th variable in our $(t,m,d)$-net to select $\ell(i)$ for any $j$. Stratification implies the ‘projection regularity’ property of Sloan and Joe (1994) in which we never get $x_{ij}=x_{i^{\prime}j}$ for any $i\neq i^{\prime}$ and any $j$.
The Sobol’ points we use are ones that have been scrambled. The scrambles we consider are either nested uniform scrambles (Owen, 1995) or random linear scrambles with digital shifts (Matoušek, 1998). Scrambling makes each $\bm{z}_{i}\sim\mathbf{U}[0,1]^{d}$ individually while $\bm{z}_{1},\dots,\bm{z}_{n}$ remain a $(t,m,d)$-net collectively.
We will estimate $\mu$ by $\hat{\mu}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})$ where $\bm{z}_{1},\dots,\bm{z}_{n}$ are a scrambled $(t,m,s+1)$-net in base $2$. This scrambled net integration has several desirable properties listed in Owen and Rudolf (2020):
@@
The scrambled Sobol’ nets are stratified.
From point 3 above we know that ${\mathrm{Var}}(\hat{\mu})=o(n^{-1})$ for $f\in L_{2}$. The Sobol’ nets have $D_{n}^{*}=O(n^{-1}\log(n)^{d-1})$ (Niederreiter, 1992, Theorem 4.10). Scrambled nets are still nets, so this holds for them too. Then by the Koksma-Hlawka inequality, ${\mathrm{Var}}(\hat{\mu})=O(n^{-2+\epsilon})$ holds for any $\epsilon>0$ when $V_{{\mathrm{HK}}}(f)<\infty$, which we describe as $f$ being in BVHK. For regular $f$ described in point 4 above we have ${\mathrm{Var}}(\hat{\mu})=O(n^{-3+\epsilon})$. Higher order digital nets can attain even better convergence rates under correspondingly stronger regularity (Dick, 2011), though they are not widely used for higher dimensions. It is common to write these powers of $\log(n)$ as $O(n^{\epsilon})$. For a discussion of that point, see Owen and Pan (2022).
While convergence $O(n^{-\rho})$ for $\rho\in\{1,2,3\}$ are widely studied, these come from error bounds and the actual convergence rate need not be near an integer. Liu (2025) shows examples where certain singular behavior of the integrand can give any rate between $1$ and $2$. In applications, the precise rate of convergence might not be known before computation begins.

### v3

@@ 1 Introduction
We study randomized quasi-Monte Carlo (RQMC) integration in a setting where one of the input variables is categorical, taking only some finite number $L\geqslant 2$ of different values. The main motivation is mixture sampling where the variables of interest come from $L$ different distributions representing $L$ different scenarios. That is, $\bm{x}\sim P_{\ell}$ with probability $\alpha_{\ell}>0$ where $\sum_{\ell=1}^{L}\alpha_{\ell}=1$. We want to estimate $\mu={\mathbb{E}}(g(\bm{x}))=\sum_{\ell=1}^{L}\alpha_{\ell}{\mathbb{E}}(g(\bm{x})\!\mid\!\bm{x}\sim P_{\ell})$.
A straightforward approach to this problem is to simply replace a standard Monte Carlo (MC) workflow with RQMC sampling. In that workflow, one random variable selects component $\ell$ with probability $\alpha_{\ell}$. Then $s$ more random variables generate $\bm{x}\sim P_{\ell}$. RQMC estimates over $s+1$ dimensions converge at a better rate than MC estimates do. Then the optimal sampling allocations become more nearly equal than proportional sampling, and we modify RQMC to take advantage of this phenomenon. We see reduced variance with those choices, and under regularity conditions, non-proportional sampling of the $P_{\ell}$ can even improve the convergence rate. We also show a minimaxity property for equal or nearly equal sampling fractions.
The input points leading to $\bm{x}\sim P_{\ell}$ come from $L$ disjoint subsets or strata in $0,1]^{s+1}$. From here on we refer to the mixture components as strata to avoid confusion with vector components.
@@
An alternative to the standard MC workflow is to compute $L$ different independently sampled RQMC estimates, one per stratum, and combine them according to their mixture probabilities $\alpha_{\ell}$. That approach misses a potential variance reduction from RQMC sampling when the within-stratum estimates are negatively correlated. Such negative correlation need not happen; we give a reason to expect them at the end of the paper.
@@
RQMC does best under favorable regularity assumptions, such as continuity, on the integrand. The presence of the categorical variable makes the resulting integrand discontinuous. That discontinuity is however axis-parallel and so it is a QMC-friendly discontinuity in the terminology of Wang & Tan (2013). In some very regular cases, RQMC can attain a mean squared error (MSE) of $O(n^{-3+\epsilon})$. Such a rate can only be attained over a sequence of special sample sizes that grow at least exponentially (Sobol’ 1998, Owen 2016). This rate can be attained by scrambled Sobol’ points with sample size equal to a power of $2$. Even an integrand with a QMC-friendly axis-parallel discrepancy is not generally integrated with an error of this rate. By using an importance sampling (Owen 2013, Chapter 9) adjustment, we can arrange to obtain $n_{\ell}=2^{m_{\ell}}$ points from $P_{\ell}$ while also having $n=\sum_{\ell=1}^{L}n_{\ell}$ be a power of $2$. This can then give the $O(n_{\ell}^{-3+\epsilon})$ rate within every stratum and hence overall.
@@
This paper is organized as follows. Section 2 introduces our notation, gives some background on RQMC, and proves some results about within-stratum discrepancies. Section 3 describes how to allocate samples to strata. We see that in RQMC sampling, it is advantageous to use more nearly equal allocations than those given by the stratum sizes. Section 4 shows how to partition a sample size $n=2^{m}$ into $L$ within-stratum sample sizes $n_{\ell}=2^{m_{\ell}}$. Section 5 establishes some minimax variance results from using equal or nearly equal sample sizes for the strata. Section 6 shows a simple toy example with $L=8$ strata where RQMC with $n_{\ell}$ taken to be powers of two does best. Section 7 uses a mixture model version of the Saint-Venant flood example from Limbourg & De Rocquigny (2010) where non-nominal sampling proportions do best. The code used to produce the results for both examples is publicly available at https://github.com/hoval58/qmc4categorical. Section 8 has some discussion and conclusions. There is an appendix for the proofs.
@@
We finish this introduction by describing some related work in the literature. The key differences between our work and the prior work we found are in how the problem has been formulated, and how the values are sampled.
@@
We use the first component of our RQMC points to select the stratum to sample from. Keller & Grünschloß (2012) used that approach in parallel computation. Their first component was used to determine which processor $\ell\in\{1,2,\dots,L\}$ got the $i$’th job. The remaining components of their QMC points sampled that job on processor $\ell$. In that setting, $\alpha_{\ell}=1/L$ and it makes sense to choose $L$ to be a power of $2$ at least for Sobol’ sequences.
@@
Cui et al. (2025) consider a target distribution on a bounded hyperrectangle $\bm{a},\bm{b}]\subset\real^{d}$. The target probability density function $\pi$ is bounded. They approximate it by a sum of products of hat functions. That is equivalent to a mixture of distributions whose components have independent triangle distributions. They consider adaptive methods to select those triangle densities given an unnormalized $\bar{\pi}\propto\pi$. For a possibly large number of mixture components, they sample each component with $n_{j}\doteq n\alpha_{j}$ input values. They use the first $n_{j}$ points of the same QMC point set for each of the strata.
@@
We use different RQMC points in each stratum. Our inputs are randomized inputs, and so we expect better error cancellation than we would get using the same inputs for every stratum. Since our analysis is based on variances, we do not require a bounded domain or integrand. As mentioned above, we use non-proportional allocation.
@@
Klebanov & Sullivan (2026) describe using transport maps from a $\mathcal{N}(0,I)$ reference distribution to sample a mixture like $P=\sum_{\ell=1}^{L}\alpha_{j}\mathcal{N}(\mu_{j},\Sigma_{j})$. They construct a continuous mapping over $t\in[0,1]$ with $\mathcal{N}(0,I)$ at $t=0$ reaching $P$ at $t=1$ and solve an ordinary differential equation along that path. For $\ell<L$, they take the first $n_{\ell}=\lfloor n\alpha_{\ell}\rfloor$ points of a quasi-Monte Carlo sequence and use those to sample from component $\ell$. Then $n_{L}=n-\sum_{\ell=1}^{L-1}n_{\ell}$.
@@ 2 Background
Here we introduce notation to describe the different variables used in the integration workflow. Then we define digital nets and discuss their randomizations. For a broader survey of RQMC see L’Ecuyer & Lemieux (2002).
We use $\mathbb{N}$ for the set of positive integers. We use $\bm{1}\{A\}$ for the indicator function that equals $1$ when $A$ holds and is 0 otherwise.
@@ 2.1 Integration variables
Our goal is to estimate this weighted sum of $d$-dimensional integrals
$\mu=\sum_{\ell=1}^{L}\alpha_{\ell}\int_{\real{}^{d}}g(\bm{x})p_{\ell}(\bm{x})\mathrm{\,d}\bm{x}$
@@
[…]
@@
Because each RQMC point $\bm{z}_{i}$ is uniformly distributed on $0,1]^{s+1}$, the estimator $\hat{\mu}$ is unbiased for $\mu$.
We also define the within-stratum integrands individually as functions of $0,1]^{s}$. The within-stratum integrands and integrals are
@@
At this point we can introduce a small generalization. Since we already have a dependence on $\ell$ in $h_{\ell}$ we could also replace $h_{\ell}(\bm{u})=g(\phi_{\ell}(\bm{u}))$ by $h_{\ell}(\bm{u})=g_{\ell}(\phi_{\ell}(\bm{u}))=g_{\ell}(\bm{x})$ to allow the function being applied to $\bm{x}$ to vary with $\ell$. That is $\mu=\sum_{\ell=1}^{L}\alpha_{\ell}{\mathbb{E}}(g_{\ell}(\bm{x})\!\mid\!\bm{x}\sim P_{\ell})$.
@@ 2.2 Sampling the strata
While $v=1$ has probability zero, for completeness we set $\bm{\ell}(v)=L$ when $v=1$. Because $v\sim\mathbf{U}[0,1]$, the expected number of samples from stratum $\ell$ is $n\beta_{\ell}$. We will always order the strata so that $\beta_{1}\geqslant\beta_{2}\geqslant\cdots\geqslant\beta_{L}$. This ordering brings a sampling advantage in Section 2.5. It is ordinarily the same ordering as the $\alpha_{\ell}$, though we consider exceptions at the end of Section 3.1.
We will assume that the sampling values $v_{1},\dots,v_{n}$ that we use are ‘stratified’. By this we mean that every interval $I_{k}=[(k-1)/n,k/n)$ contains precisely one of $v_{1},\dots,v_{n}$ for $k=1,\dots,n$. The next proposition shows that $n_{\ell}>0$ whenever $\beta_{\ell}>2/n$.
@@
For $n\geqslant 1$, let $v_{1},\dots,v_{n}$ be stratified. For $A,B)\subset[0,1)$, let $n_{*}$ be the number of $v_{i}\in[A,B)$. Then
@@
[…]
@@ 2.3 Local and star discrepancy
For $\bm{a}\in[0,1)^{d}$, the anchored box $\bm{0},\bm{a})$ has volume $\mathrm{Vol}([\bm{0},\bm{a}))=\prod_{j=1}^{d}a_{j}$. The local discrepancy of $n$ points $\bm{x}_{1},\dots,\bm{x}_{n}\in[0,1)^{d}$ at $\bm{a}$ is
@@
The Koksma-Hlawka inequality (Hickernell 2014) is that
@@ 2.4 Digital nets
Here we present the defining properties of digital nets. They exist in integer bases $b\geqslant 2$ (Dick & Pillichshammer 2010) but we only consider base $2$ here. For a dimension $d\geqslant 1$, elementary intervals in base $2$ are dyadic hyperrectangles of the form
@@
for integers $k_{j}\geqslant 0$ and $0\leqslant c_{j}<2^{k_{j}}$. We use $|\bm{k}|=\sum_{j=1}^{d}k_{j}$, so $H(\bm{c},\bm{k})$ has volume $2^{-|\bm{k}|}$.
For integers $m\geqslant t\geqslant 0$ the $n=b^{m}$ points $\bm{z}_{1},\dots,\bm{z}_{n}\in[0,1)^{d}$ are a $(t,m,d)$-net in base $b$ if $H(\bm{c},\bm{k})$ contains precisely $nb^{-|\bm{k}|}=b^{m-|\bm{k}|}$ of them, whenever $m-|\bm{k}|\geqslant t$. Smaller $t$ are better. We always have $t\leqslant m$ because $t=m$ just means that all the points are in $0,1)^{d}$.
A $(t,d)$-sequence in base $2$ is an infinite sequence of values $\bm{x}_{i}\in[0,1]^{d}$, with the following property: any subsequence $\bm{x}_{a2^{r}+1},\dots,\bm{x}_{(a+1)2^{r}}$ for integers $a\geqslant 0$ and $r\geqslant m$ is a $(t,r,d)$-sequence in base $2$. With a $(t,d)$-sequence we get an infinite sequence of $(t,m,d)$-nets. Consecutive pairs of those $(t,m,d)$-nets yield an infinite sequence of $(t,m+1,d)$-nets of which pairs form $(t,m+2,d)$-nets and so on without limit. We will use digital nets in base $2$ taken from Sobol’ sequences using the direction numbers from Joe & Kuo (2008).
We call a $(t,m,d)$-net in base $2$ stratified if for each component $j=1,\dots,d$, the values $x_{1j},\dots,x_{nj}$ are stratified. A net with $t=0$ is automatically stratified, but some digital nets with $t>0$ are also stratified. The $(t,m,d)$-nets from Sobol’ sequences are stratified. The stratification property means we could use the $j$’th variable in our $(t,m,d)$-net to select $\ell(i)$ for any $j$. Stratification implies the ‘projection regularity’ property of Sloan & Joe (1994) in which we never get $x_{ij}=x_{i^{\prime}j}$ for any $i\neq i^{\prime}$ and any $j$.
The Sobol’ points we use are ones that have been scrambled. The scrambles we consider are either nested uniform scrambles (Owen 1995) or random linear scrambles with digital shifts (Matoušek 1998). Scrambling makes each $\bm{z}_{i}\sim\mathbf{U}[0,1]^{d}$ individually while $\bm{z}_{1},\dots,\bm{z}_{n}$ remain a $(t,m,d)$-net collectively.
We will estimate $\mu$ by $\hat{\mu}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})$ where $\bm{z}_{1},\dots,\bm{z}_{n}$ are a scrambled $(t,m,s+1)$-net in base $2$. This scrambled net integration has several desirable properties listed in Owen & Rudolf (2020):
@@
The scrambled Sobol’ nets are stratified. Digital nets in base $2$ have sample sizes $n=2^{m}$ by definition. Arbitrary sample sizes $n$ are available as the first $n$ points of scrambled Sobol’ sequences. Property $1$ holds for any finite $n$ there and Property 2 holds for any sequence $n\to\infty$ (Owen & Rudolf 2020). The other properties are specific to $n=2^{m}$.
From point 3 above we know that ${\mathrm{Var}}(\hat{\mu})=o(n^{-1})$ for $f\in L_{2}$. The Sobol’ nets have $D_{n}^{*}=O(n^{-1}\log(n)^{d-1})$ (Niederreiter 1992, Theorem 4.10). Scrambled nets are still nets, so this holds for them too. Then by the Koksma-Hlawka inequality, ${\mathrm{Var}}(\hat{\mu})=O(n^{-2+\epsilon})$ holds for any $\epsilon>0$ when $V_{{\mathrm{HK}}}(f)<\infty$. Then $f$ is of bounded variation in the sense of Hardy and Krause, which we describe as $f$ being in BVHK. For regular $f$ described in point 4 above we have ${\mathrm{Var}}(\hat{\mu})=O(n^{-3+\epsilon})$. Higher order digital nets can attain even better convergence rates under correspondingly stronger regularity (Dick 2011), though they are not widely used for higher dimensions. It is common to write these powers of $\log(n)$ as $O(n^{\epsilon})$. For a discussion of that point, see Owen & Pan (2022).
While variance rates $O(n^{-\rho})$ for $\rho\in\{1,2,3\}$ are widely studied, these come from error bounds. The actual convergence rate need not be near an integer. Liu (2025) shows examples where certain singular behavior of the integrand can give any rate between $1$ and $2$. In applications, the precise rate of convergence might not be known before computation begins.